OpenStudy (anonymous):

A customer from Cavallaro's Fruit Stand picks a sample of 5 oranges at random from a crate containing 65 oranges, of which 6 are rotten. What is the probability that the sample contains 1 or more rotten oranges? (Round your answer to three decimal places.

OpenStudy (kropot72):

P(zero rotten) + P(1 or more rotten) = 1 Therefore if we find the probability of no rotten fruit in the sample, and subtract that probability from 1 we will have the required probability. Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Good. \[P(0\ rotten)=\frac{C(6, 0) \times C(59, 5)}{C(65, 5)}=you\ can\ calculate\]

OpenStudy (anonymous):

1.6498

OpenStudy (kropot72):

Not really. For a start the value of probability can never be more than 1. Are you using a calculator to find the value?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

0.606

OpenStudy (kropot72):

Right. Can you tell me the value of: C(6, 0) * C(59, 5) = ?

OpenStudy (anonymous):

1*5006386

OpenStudy (kropot72):

Correct! No the value of: C(65, 5) = ?

OpenStudy (anonymous):

8259888

OpenStudy (kropot72):

Correct again! So \[P(0\ rotten)=\frac{5006386}{8259888}=you\ can\ calculate\]

OpenStudy (anonymous):

0.606

OpenStudy (kropot72):

Correct! Therefore we get: P(one or more rotten) = 1.000 - 0.606 = ?

OpenStudy (anonymous):

0.394

OpenStudy (kropot72):

That is what I get also.

OpenStudy (anonymous):

thank you : )

OpenStudy (kropot72):

You're welcome :)